Could you please give step by step solutions and explain how you get the final answer?

Duration of bond ABC is 6.67 years and its convexity is 135. If that bond has current price of £107, yield to maturity of 5% and if yields decrease by 1.25%, what would be the new price of this bond? Explain.
Answer #1

Duration = 6.67 years

Convexity = 135

Change in Yield = ΔYield = -1.25% = -0.0125

Hence, Change in Price = (−Duration * ΔYield)+(1/2 * Convexity * (ΔYield)2)

=  (−6.67 * (-0.0125))+(1/2 * 135 * (-0.0125)2) = 0.0939 or 9.39%

Hence, Price of bond increases by 9.39%

New Price of bond = Current Price of Bond * (1 + change in price) = 107(1+0.0939) = £117.05